OpenStudy (anonymous):

Which correctly describes how the graph of the inequality 5x + 2y ≥ 13 is shaded? Above the solid line Below the solid line Above the dashed line Below the dashed line

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What do you get when you isolate y

OpenStudy (anonymous):

I'm not sure, can you explain how to do that?

jimthompson5910 (jim_thompson5910):

5x + 2y is the same as 2y + 5x

jimthompson5910 (jim_thompson5910):

what will undo the +5x ?

OpenStudy (anonymous):

adding or subtracting 5x to both sides?

jimthompson5910 (jim_thompson5910):

which one?

OpenStudy (anonymous):

subtracting?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so \[\Large 2y \ge 13 - 2x\] \[\Large 2y \ge -2x + 13\]

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

now you subtract 2x from both sides?

jimthompson5910 (jim_thompson5910):

we want to isolate y

OpenStudy (anonymous):

divide by 2?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

\[\Large 2y \ge -2x + 13\] \[\Large y \ge \frac{-2x + 13}{2}\]

jimthompson5910 (jim_thompson5910):

is the boundary solid? or dashed?

OpenStudy (anonymous):

solid?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

where is the shading

OpenStudy (anonymous):

below the line

jimthompson5910 (jim_thompson5910):

why below?

OpenStudy (anonymous):

oh sorry I meant above

jimthompson5910 (jim_thompson5910):

why above?

OpenStudy (anonymous):

Because it is greater than or equal to

jimthompson5910 (jim_thompson5910):

correct, the y coordinates of the points in the solution set will be above the line

OpenStudy (anonymous):

Yay!!! ^-^

OpenStudy (anonymous):

Thanks!